OpenStudy (anonymous):

If you have fractions of 1/2 and 19/2 which of the two are f(x) is it the x intercept? Does the graph open up or down?

OpenStudy (anonymous):

wish i could understand mind explaining more

OpenStudy (anonymous):

i'm lost - cant understand a word!

OpenStudy (anonymous):

I have figured out the x intercept is 1/2 and y intercept is 19/2 and I have to figure out f(x) which I think I figured out the answer to f(x) is always y, I think. But I am not sure in graphing if it opens up or down?

OpenStudy (anonymous):

then 1/2 is your gradient right? have you tried sketching?

OpenStudy (anonymous):

yes, I think I have got it, thanks anyway?

OpenStudy (anonymous):

Say you have an equation, I dunno, y = 2x + 9. You get the y-intercept by putting x = 0 so y is 9 and (0,9) is the intercept. And the x-intercept by putting y = 0 so x is 4 and a half, (0, 9/2) is the intercept. Instead of writing y = 2x + 9 I can write f(x) = 2x + 9 which means of course, that y= f(x). If I say make the graph of f(x) = 2x + 9 or make the graph of y = 2x + 9 it's the exact same graph except instead of writing y on the "up" axis, I write f(x) instead.